Please help me on this question. In 1990, the population of Alabama was 4,040,587. In 2000, the population was 4,447,100. Find the percent of change from 1990 to 2000. i know the answer is 10.1% I just dont know how to work it. please help. thank you.
Answer by cdafonseca(2)

4447100-4040587=406513 (406513/4040587)*100=10.060...%

~10.1%